OpenStudy (anonymous):

In a right triangle, one of the acute angles is five times as large as the other acute angle. Find the measure of the two acute angles.

OpenStudy (kc_kennylau):

Let the smaller angle be \(\theta\), then the bigger angle would be \(5\theta\)

OpenStudy (kc_kennylau):

We have \(\theta+5\theta+90^\circ=180^\circ\) get it?

OpenStudy (anonymous):

Yes, thank you so much!

OpenStudy (anonymous):

The answers were 15 and 75.